The cost of renting a banquet hall for an event is $300 plus $30 for each person attending the event. If the hall provides live music, the cost is 40% more per person.
A). How much is the cost of an event for 125 people with like music ?
B). write a function for the cost of an event that includes luve music.

thanks

Found 2 solutions by Boreal, rfer:
Answer by Boreal(15235)

C=$300+$30P
C=$300+$30*1.4P=$300+$42P--live music function
C=$300+$42*125=$5550

Answer by rfer(16322)

300+(30*1.40)(125)=$5550